1. The area of a rectangle is 60 square meters, and the width is 6 meters. find the length.
2. The length of a rectangle is 15 centimeters, and the perimeter is 50 centimeters. Find the width.

Answer by MArk_HeRras(6)

In Question No.1
A/Representing: X =Length
6 =Width
Equation:2(6) + 2(x) = 60
Solution:
2(6) + 2(x) = 60
12 + 2x = 60
2x = 60 - 12
2x = 48
2x/2 = 48/2
x = 24
now for checking
A: x - 24 =length
6 =width
Checking:2(6) + 2(24) = 60
12 + 48 = 60
60 =60


And last quesrion No.2
A/representing: x =width
15 =length
E: 2(x) + 2(15) = 50
S:
2(x) + 2(15) = 50
2x + 30 = 50
2x=50-30
2x/2 = 20/2
x = 10
Now checking
A: x = 10 = width
15 = length
C: 2(10) + 2(15) = 50
20 + 30 = 50
50 = 50


So the answer in No.1 is x - 24 =length,6 =width
And in No.2 is x = 10 = width,15 = length